# Unleashing the Potential of Large Language Models as Prompt Optimizers: An Analogical Analysis with Gradient-based Model Optimizers

### v2

@@ Abstract
Automatic prompt optimization is an important approach to improving the performance of large language models (LLMs). Recent research demonstrates the potential of using LLMs as prompt optimizers, which can generate improved task prompts via iterative refinement. In this paper, we propose a novel perspective to investigate the design of LLM-based prompt optimizers, by drawing an analogy with gradient-based model optimizers. To connect these two approaches, we identify two pivotal factors in model parameter learning: update direction and update method. Focused on the two aspects, we borrow the theoretical framework and learning methods from gradient-based optimization to design improved strategies for LLM-based prompt optimizers. By systematically analyzing a rich set of improvement strategies, we further develop a capable Gradient-inspired LLM-based Prompt Optimizer called GPO. At each step, it first retrieves relevant prompts from the optimization trajectory as the update direction. Then, it utilizes the generation-based refinement strategy to perform the update, while controlling the edit distance through a cosine-based decay strategy. Extensive experiments demonstrate the effectiveness and efficiency of GPO. In particular, GPO brings an additional improvement of up to 56.8% on Big-Bench Hard and 55.3% on MMLU compared to baseline methods. The code is available at https://github.com/RUCAIBox/GPO.
Unleashing the Potential of Large Language Models as Prompt Optimizers: An Analogical Analysis with Gradient-based Model Optimizers
Xinyu Tang1,3thanks:   Equal contribution., Xiaolei Wang1,311footnotemark: 1, Wayne Xin Zhao1,3thanks:   Corresponding author., Siyuan Lu, Yaliang Liand Ji-Rong Wen1,2,3 1Gaoling School of Artificial Intelligence, Renmin University of China 2School of Information, Renmin University of China 3Beijing Key Laboratory of Big Data Management and Analysis Methods 4International School, Beijing University of Posts and Telecommunications 5Alibaba Group txy20010310@163.com, wxl1999@foxmail.com, batmanfly@gmail.com
1 Introduction
@@
Update direction Model value Task prompt
Gradient Reflection
@@
Figure 1: The overview of LLM-based prompt optimization framework.
Nowadays, prompting has become the pivotal approach to unleashing the power of large language models (LLMs) Touvron et al. (2023); OpenAI (2023). However, prompt engineering is not easy and requires extensive trial-and-error efforts since LLMs are sensitive to prompts Zhao et al. (2021); Lu et al. (2022); Wei et al. (2023). Although general guidelines for high-quality prompts exist Kojima et al. (2022); Amatriain (2024), they cannot always lead to optimal task performance.
To improve the task performance of LLMs, automatic prompt optimization has been proposed Zhou et al. (2023). Early work either directly performs discrete optimization through methods like reinforcement learning Deng et al. (2022); Zhang et al. (2023) or performs continuous optimization in the embedding space of LLMs Shin et al. (2020); Lin et al. (2023); Chen et al. (2023). However, these methods often require access to the logits or internal states of LLMs, which is infeasible for those only accessible through APIs. In addition, they need to be specially trained for each task. Considering these issues, recent work proposes to model the optimization problem in natural language and using LLMs as prompt optimizers Zhou et al. (2023); Ma et al. (2024); Schnabel and Neville (2024). In this approach, LLMs can perform optimization with only outputs from the task model and quickly adapt to various tasks without training. Figure 1 illustrates the overall framework of LLM-based prompt optimization. However, such a method raises a new challenge for the design of meta-prompt, which is the prompt for LLMs to perform prompt optimization.
To tackle this issue, we aim to investigate the design of meta-prompts. In existing approaches, meta-prompts are often hand-crafted Yang et al. (2023) or optimized with heuristic algorithms Fernando et al. (2023). Despite the flexibility, these studies still lack principled guidelines about their designs. Our work is inspired by the great success of gradient-based optimizers in model optimization, which have been systemically studied in both theory and practice Sun et al. (2020); Abdulkadirov et al. (2023). Since both optimizers target enhancing model performance through optimization, it is feasible to connect the two different approaches via analogical analysis. In this way, we can borrow the theoretical framework and extensive research of gradient-based model optimizers to enhance LLM-based prompt optimizers.
Specifically, in this paper, we take the gradient-based optimization approach from machine learning as reference, and identify two key factors for model parameter learning, namely update direction (i.e., gradient calculation) and update method (i.e., gradient descent), as presented in Table 1. By analogy, in LLM-based prompt optimizers, the update direction can refer to suitable prompt information informing the LLM about how to produce more effective task prompts, while the update method can refer to how the LLM utilizes such information to improve prompts. By drawing such an analogy, we can develop a more formal framework for LLM-based prompt optimizers and provide guidelines to design more principled meta-prompts.
Furthermore, we develop a capable Gradient-inspired LLM-based Prompt Optimizer called GPO, with the best configuration from our systematic study. We evaluate its performance across complex reasoning, knowledge-intensive, and common NLP tasks in various evaluation settings. When using Llama-2-7b-chat as the task model, the prompts produced by GPO surpass the instruction “Let’s think step by step” by 18.5% on Big-Bench Hard (BBH) and 7.6% on MMLU. Furthermore, GPO produces an additional improvement of up to 56.8% on BBH and 55.3% on MMLU compared with baseline methods while using fewer tokens.
To the best of our knowledge, it is the first time that a systematic study has been conducted for LLM-based prompt optimizers. More specifically, it has been studied by analogy with gradient-based model optimizers, which we believe is useful to seek theoretical foundations and extend feasible approaches for prompt optimization.
2 An Analogical Analysis Between Gradient-Based Model Optimizer and LLM-Based Prompt Optimizer
In this section, we present an analogical analysis between model optimization and prompt optimization to build their connections and further improve existing LLM-based prompt optimizers.
2.1 Task Formulation
Prompt optimization aims to find the optimal task prompt $p^{*}$ in the format of natural language that maximizes the performance on a specific task dataset $\mathcal{D}$ when using an LLM as the task model $\mathcal{M}_{T}$. To perform such optimization, our idea is to develop a prompt optimizer, which can be built upon some search algorithm (e.g., evolutionary algorithms Guo et al. (2023)) or an LLM Yang et al. (2023). In this paper, we focus on using an LLM as the prompt optimizer $\mathcal{M}_{O}$. Formally, the problem of prompt optimization can be formulated as:
@@ 2.1 Task Formulation
where $p$ is the prompt generated by the LLM-based prompt optimizer $\mathcal{M}_{O}$, $\mathcal{M}_{T}(x;p)$ represents the output from the task model for input $x$ conditioned on the prompt $p$, and the function $F(\cdot)$ calculates the task performance based on some measurement.
For the LLM-based prompt optimizer, it requires another prompt to perform the optimization for the task prompt, which is usually called meta-prompt Yang et al. (2023); Ye et al. (2023). For example, for the mathematical reasoning task, the prompt can be “Let’s solve the problem”, while the meta-prompt may be “Improve the prompt to help a model better perform mathematical reasoning”.
2.2 Analogical Analysis for Prompt Optimization
Since both prompt optimizers and model optimizers target enhancing the model performance through some optimization algorithm, in this part, we aim to draw inspiration from the design of gradient-based model optimizers to conduct a systematic analysis of LLM-based prompt optimizers.
2.2.1 Revisiting Gradient-Based Optimizers
Similar to prompt optimization, model optimization aims to find the optimal values of model parameters that minimize the loss function. In model optimization, gradient-based optimizers are the most widely used approaches, which iteratively update model parameters in the direction of the negative gradient of the loss function. To motivate our approach, we take the fundamental optimizer known as gradient descent Boyd and Vandenberghe (2014) for discussion.
In the basic form of gradient descent, a single optimization step can be formulated as follows:
@@ 2.2.1 Revisiting Gradient-Based Optimizers
where $\Theta_{k}$ and $\Theta_{k+1}$ are the values of model parameters at the last and current steps, $\tau_{k}$ and $g_{k}$ are the learning rate and gradient at the current step. Gradient descent can be improved by focusing on two elements in the formula: $\tau_{k}$ and $g_{k}$. For $\tau_{k}$, learning rate schedulers Gotmare et al. (2019) are proposed to dynamically adjust the learning rate. For $g_{k}$, the concept of momentum Sutskever et al. (2013) is introduced to include historical gradients, and its computation can be expressed as follows: $v_{k+1}=\beta v_{k}+g_{k}=\sum_{i=1}^{k}\beta^{k-i}g_{i}$, where $\beta$ represents the momentum coefficient.
@@
Despite various gradient-based optimizers, they mainly model two key factors, namely update direction (e.g., gradient $g_{k}$) and update method (e.g., direct descent update by subtracting $\tau_{k}g_{k}$). Our approach is inspired by the observation that existing LLM-based prompt optimization methods also implicitly employ the two aspects. For example, OPRO Yang et al. (2023) uses previous task prompts along with their performance to guide the direction of update and utilizes in-context learning based generation as the update method. The comparison between glossaries is shown in Table 1. However, the efforts in existing work only initially explore the design of the two key factors, and we aim to conduct more in-depth and systematic investigations by borrowing the idea of research in gradient-based optimization.
2.2.2 Analogical Prompt Optimization Strategies for “Update Direction
In prompt optimization, there are no explicit gradients for controlling the update direction. However, we can incorporate the concept of gradient into the design of meta-prompts to improve the prompt optimization process.
Analogical “Gradient” Forms. The basic function of the gradient is to inform how the optimization process should adjust according to the model performance. To mimic similar effects, for an LLM with fixed parameters, it is only feasible to design suitable prompting strategies to adjust the output. Here, we consider two analogical forms to implicitly support the gradient-like function.
$\bullet$ Prompt+performance. One straightforward method is to include the last-round task prompt and the corresponding model performance into the meta-prompt for LLM-based optimizer $\mathcal{M}_{O}$. It leverages the capacity of LLMs to reason about how to improve prompting optimization.
$\bullet$ Prompt+performance+reflection. Another way to solve the barrier of the gradient is to leverage the reflection capability of LLMs Pryzant et al. (2023). With the reflection mechanism, LLMs can generate feedback from past failures, which can be used to improve performance. Such feedback can be seen as a form of “semantic” gradient signals Pryzant et al. (2023).
Analogical “Momentum” Forms. Inspired by the momentum method Sutskever et al. (2013) in gradient descent, we consider enhancing the aforementioned basic form of meta-prompt by leveraging the intermediate results accumulated in the prompt optimization process. A straightforward way is to directly include them into the meta-prompt. However, it might be limited by the context length of LLMs and affected by the accumulated noise. To better utilize the optimization trajectory, we propose two alternative methods.
$\bullet$ Summarization-based trajectory. One direct approach is to summarize the intermediate results from the optimization trajectory. Specifically, at each step, we use an instruction (e.g., “Your task is to summarize the problems in the previous prompt and the current prompt.”) to let the LLM perform summarization using the summary in the last step and the result in the current step.
$\bullet$ Retrieval-based trajectory. Another approach is to dynamically retrieve $k$ pieces of gradients from the optimization trajectory. Specifically, we consider the following three strategies for selection: (1) recency: selecting $k$ nearest gradients; (2) relevance: selecting $k$ most relevant gradients, which are measured by the semantic similarity based on the BGE model Xiao et al. (2023); (3) importance: selecting $k$ most important gradients, which is measured by the performance gain.
2.2.3 Analogical Prompt Optimization Strategies for “Update Method
Another important factor to consider is the update method for prompt optimization. In gradient-based model optimizers, the optimization process is controlled by the learning rate and fulfilled by gradient descent on parameters. Accordingly, we explore two analogical aspects: prompt variation control and meta-prompt refinement.
Prompt Variation Control. After setting the optimization direction, it is important to adjust the parameter optimization process with a suitable learning rate. This is because using a either too large or too small rate can result in an oscillating or slowly converging optimization process. Similarly, without a suitable control mechanism in prompt optimization, the LLM-based optimizer might overshoot the optimal prompt or oscillate in the optimization process. To mimic the similar effects of learning rate, our idea is to control the variation degree of prompt optimization, which is measured by the edit distance between two task prompts at consecutive iterations. Specifically, following Ye et al. (2023), we add an instruction (i.e., “You are allowed to change up to {number} words in the current prompt.”) into the meta-prompt to limit the number of words that can be modified. Accordingly, we propose two control methods as follows:
$\bullet$ Decay-based constraint. To avoid overshooting the optimal solution, the decay strategy is proposed to gradually reduce the learning rate Izmailov et al. (2018). Here, we borrow the idea to control the maximum edit distance and consider gradually reducing its value following either a linear or cosine curve. In addition, following the common practice in training LLMs Touvron et al. (2023), we reduce the constraint to approximately 20% of its maximum value until convergence.
$\bullet$ Warmup-based constraint. To avoid instability in the early stage of optimization, the warmup strategy is proposed to gradually increase the learning rate at the beginning Goyal et al. (2017). Similarly, we adopt the widely used linear warmup schedule to gradually increase the constraint for the maximum edit distance to its initially set value in the initial 5% steps.
Prompt Refinement Strategy. During the optimization process, the task prompt would be iteratively refined, to improve the corresponding model performance. By analogy with the subtraction operation in gradient-based optimizers (i.e., $-\tau_{k}g_{k}$ in Eq. (2)), we introduce two methods to update the task prompt accordingly.
$\bullet$ Editing-based refinement. The first method directly edits the last-round task prompt to improve performance Pryzant et al. (2023). Specifically, we add an instruction (i.e., “Modify the current prompt and get a new improved promp”) into the meta-prompt, which requires the LLM to edit the task prompt from the last step according to the update direction. This method allows for effectively exploiting the existing prompt, leading to a gradual performance improvement through a relatively stable optimization process.
$\bullet$ Generation-based refinement. In addition to direct edits, we can also leverage the in-context learning capability of LLMs to generate refined task prompts Zhou et al. (2023). Specifically, we present the information regarding the update direction in the format of demonstration. Then, the LLM can follow the demonstration to generate a new task prompt. Compared with the editing-based updating method, the generation-based approach explores a wider range of prompt variations, which has the potential to yield improved performance but also makes the optimization process unstable.
2.3 Analogical Analysis Experiments
In this part, we first describe the experiment setting for our analogical analysis, and then detail our findings from the experiment.
2.3.1 Experiment Setup
We select a dataset from each type of task in BBH Suzgun et al. (2023) to create a lite BBH benchmark for our analysis: i) Navigate (binary choice); ii) Movie Recommendation (multiple choice); iii) Object Counting (numeric response); iv) Word Sorting (free response). We adopt exact match as the metric for performance evaluation. We employ Llama-2-7b-chat Touvron et al. (2023) as the task model and gpt-3.5-turbo as the prompt optimizer. The initial task prompt is “Let’s think step by step.” Other details are presented in Appendix C.
2.3.2 Empirical Findings
Initial prompt: “Let’s think step by step.” Performance: 31.25
Gradient form Momentum form
None S-T
R-T
Recency Relevance Importance
P+M
41.07 41.03 41.93 42.53 41.84
P+M+R
40.34 40.63 41.55 40.81 39.73
@@ 2.3.2 Empirical Findings
Table 2: The performance comparison between different update directions. “P” denotes prompt, “M” denotes performance, “R” denotes reflection, “S-T” denotes summarization-based trajectory, and “R-T” denotes retrieval-based trajectory.
Findings For Update Direction. The results for the update direction of prompt optimization are presented in Table 2. Here are the main findings:
$\bullet$ With regard to the form of gradient, prompt+ performance achieves better performance than prompt+performance+reflection, which brings an improvement of up to 31% compared with the initial task prompt. The substantial improvement brought by prompts can be attributed to their rich semantic information about the task, which can activate the task-specific knowledge of LLMs for optimization. In contrast, LLMs are known to be limited in their capabilities of reflection Huang et al. (2023), which may lead to inaccurate updates.
$\bullet$ The analogical concept of momentum can further improve the performance of prompt optimization. Among various designs, relevance-based trajectory emerges as the most effective one, which brings an additional 15% improvement. This can be attributed to the fact that LLMs can learn more from prompts that are contextually relevant, while it might be challenging for LLMs to fully understand the signal of recency or importance. By contrast, the summarization-based trajectory proves to be less helpful. This is because summarization only captures common aspects of the trajectory while neglecting details that may be crucial.
Initial prompt: “Let’s think step by step.” Performance: 31.25
Refinement strategy Editing Generation
Variation Control No control 42.53 42.61
Fixed 42.91 43.09
+Warmup 41.76 40.08
Linear decay 42.68 42.86
+Warmup 41.47 41.12
Cosine decay 42.95 43.75
+Warmup 40.19 41.29
@@
Table 3: The performance comparison between different update methods.
@@
Findings For Update Method. To investigate the update method for prompt optimization, we conduct experiments using the best configuration found in the previous experiments. The results for the update method of prompt optimization are presented in Table 3. Here are the key findings:
@@
$\bullet$ In general, generation-based refinement performs better than editing-based refinement, which brings an improvement of up to 36% compared with the initial task prompt. This can be partly attributed to that exploration is more important in prompt optimization. The generation-based strategy is not confined to the current prompt, allowing LLMs to explore a boarder search space, thereby demonstrating greater flexibility and creativity.
@@
$\bullet$ Among various controlling methods for prompt variation, cosine decay-based constraint achieves the best performance, bringing an additional 10% improvement. However, unlike gradient-based model optimization, the warmup strategy does not yield improvement. This finding suggests that, at the early stage of prompt optimization, exploration plays a crucial role, while stability becomes more important in the later stage.
3 Our Prompt Optimizer: GPO
In this section, we present our novel gradient-inspired LLM-based prompt optimizer called GPO, which leverages the insights gained from our systematic study. Our approach proposes a novel design of the meta-prompt, aiming to unleash the potential of LLMs as prompt optimizers.
Iterative Prompt Optimization. GPO performs prompt optimization through a multi-step iterative process. At each step, the LLM first generates multiple candidate task prompts based on a meta-prompt. The meta-prompt serves as the input that guides the LLM in optimizing its prompts. Subsequently, the task prompt with the best performance is selected for the next iteration. This iterative process continues until either the performance of the task prompt reaches a plateau or the predefined maximum number of optimization steps is reached.
The Design of Meta-Prompt. As the input to the LLM, our meta-prompt consists of two key components: update direction and update method.
$\bullet$ Update direction. To determine the update direction, we leverage the retrieval-based optimization trajectory in Section 2.2.2. This trajectory comprises a collection of past task prompts, along with their model performance. They are selected using a relevance-based strategy and are sorted in ascending order based on their similarity to the current prompt.
$\bullet$ Update method. After the update direction is determined, we further employ the generation-based refinement strategy in Section 2.2.3 to update the task prompt. Specifically, we present the trajectory in the format of demonstration in the meta-prompt. Then, the LLM can follow these demonstrations to generate a new task prompt via in-context learning. Additionally, we implement the cosine-based decay strategy in Section 2.2.3 with an instruction to control the edit distance between task prompts at consecutive iterations, ensuring gradual and controllable changes.
Furthermore, we enhance the meta-prompt by incorporating a few task examples. These examples provide additional context to aid the LLM in understanding the task more effectively. The complete meta-prompt is presented in Appendix D.2.

### v3

# Unleashing the Potential of Large Language Models as Prompt Optimizers: Analogical Analysis with Gradient-based Model Optimizers

@@ Abstract
Automatic prompt optimization is an important approach to improving the performance of large language models (LLMs). Recent research demonstrates the potential of using LLMs as prompt optimizers, which can generate improved task prompts via iterative refinement. In this paper, we propose a novel perspective to investigate the design of LLM-based prompt optimizers, by drawing an analogy with gradient-based model optimizers. To connect these two approaches, we identify two pivotal factors in model parameter learning: update direction and update method. By systematically analyzing a rich set of improvement strategies on the two aspects, we further develop a capable Gradient-inspired LLM-based Prompt Optimizer called GPO. At each step, it first retrieves relevant prompts from the optimization trajectory as the update direction. Then, it utilizes the generation-based refinement strategy to perform the update, while controlling the edit distance through a cosine-based decay strategy. Extensive experiments demonstrate the effectiveness and efficiency of GPO. In particular, GPO brings an additional improvement of up to 56.8% on Big-Bench Hard and 62.6% on MMLU compared to baseline methods. The code is available at https://github.com/RUCAIBox/GPO.
Introduction
Nowadays, prompting has become the pivotal approach to unleashing the power of large language models (LLMs) (Zhao et al. 2023). However, prompt engineering is not easy and requires extensive trial-and-error efforts since LLMs are sensitive to prompts (Lu et al. 2022; Wang et al. 2024; Tang et al. 2024). Although general guidelines for high-quality prompts exist (Kojima et al. 2022; Amatriain 2024), they cannot always lead to optimal task performance.
To improve the task performance of LLMs, automatic prompt optimization has been proposed (Zhou et al. 2023). Early work either performs discrete optimization through methods like reinforcement learning (Zhang et al. 2023) or performs continuous optimization in the embedding space (Chen et al. 2023; Wang et al. 2022a, 2023). However, they often require access to the logits or internal states of LLMs, which is infeasible for those only accessible through APIs. In addition, they need to be specially trained for each task. Considering these issues, recent work proposes to model the optimization problem in natural language and using LLMs as prompt optimizers due to their strong understanding (Zhan et al. 2024) and generation capabilities (Schnabel and Neville 2024). In this approach, LLMs perform optimization with only outputs from the task model and quickly adapt to various tasks without training. However, such a method raises a new challenge for the design of meta-prompt, which is the prompt for LLMs to perform prompt optimization.
Figure 1: Comparisons of GPO to existing LLM-based prompt optimizers in terms of effectiveness (Accuracy) and efficiency (improvement per dollar spent on API) on BBH.
Figure 2: The overview of the GPO framework. “Current information”, “Trajectory”, “Edit distance”, and “Refinement strategy” are concepts of LLM-based prompt optimizers, which can correspond to “Descent direction”, “Momentum”, “Learning rate”, and “Descent” in gradient-based model optimizers.
To tackle this issue, we aim to investigate the design of meta-prompts. Existing methods for creating meta-prompts typically involve either manual human effort (Yang et al. 2023) or heuristic algorithms (Fernando et al. 2023). Despite the flexibility, these studies still lack principled guidelines about their designs. Our work is inspired by the great success of gradient-based optimizers in model optimization, which have been systemically studied in both theory and practice (Sun et al. 2020). Since both optimizers aim at enhancing model performance through iterative optimization, it is feasible to connect the two different approaches via analogical analysis. In this way, we can borrow the theoretical framework and extensive research of gradient-based model optimizers to enhance LLM-based prompt optimizers.
Therefore, in this paper, we propose a comprehensive analogy framework for the two key factors (i.e., update direction and update method) in LLM-based optimizers. As illustrated in Table 1, for update direction, we analogize descent direction to the information of the current prompt and momentum to the information of the historical prompt, while for the update method, we analogize the learning rate to edit distance and gradient descent to the refinement strategy. Based on such a framework, we conduct systematic empirical studies on various implementations for each component and report their experimental results with detailed analysis.
Based on the findings from our systematic analysis, we further develop a capable Gradient-inspired LLM-based Prompt Optimizer called GPO, with the best implementation for each component. Figure 2 illustrates the overall framework of GPO. We evaluate its effectiveness across various LLMs, tasks, and evaluation settings. When using Llama-2-7b-chat as the task model, the prompts produced by GPO surpass the instruction “Let’s think step by step” by 18.5% on Big-Bench Hard (BBH) and 7.6% on MMLU. Furthermore, GPO produces an additional improvement of up to 56.8% on BBH and 62.6% on MMLU compared with baseline methods while using fewer tokens.
Our contributions can be summarized as follows:
$\bullet$ To the best of our knowledge, this is the first time that a systematic analogy has been conducted for LLM-based prompt optimizers with gradient-based model optimizers.
$\bullet$ We conduct a comprehensive experimental analysis on the two key factors (i.e., update direction and update method) and report several key findings.
$\bullet$ Based on the findings of the systematic analysis, we develop a more effective and efficient LLM-based prompt optimizer, GPO, which surpasses competitive baseline methods across various LLMs, tasks, and evaluation settings while incurring lower costs.
Related Work
Prompt Engineering and Optimization.
Prompt engineering aims to find suitable prompts for LLMs to perform various tasks. To reduce human efforts, researchers have explored automatic prompt optimization, which can be categorized into continuous and discrete optimization methods. Discrete methods directly optimize the natural language prompts through methods like reinforcement learning (Deng et al. 2022; Zhang et al. 2023) and editing (Xu et al. 2022; Prasad et al. 2023). In contrast, continuous methods perform optimization in the embedding space of LLMs, allowing for optimization through gradient (Li and Liang 2021). We focus on discrete methods, especially LLM-based prompt optimizers.
LLM-based Prompt Optimizers.
Due to the unprecedented capabilities of LLMs, recent work starts to utilize them as prompt optimizers. One line of work (Guo et al. 2023; Yang and Li 2023) combines LLMs with evolutionary algorithms to perform prompt optimization. Another line of work (Yang et al. 2023) aims to adapt concepts and techniques from gradient-based model optimizers (e.g., gradient (Pryzant et al. 2023) and momentum (Yang et al. 2023)) to LLM-based prompt optimizers. However, no comprehensive guidelines exist for using LLMs as prompt optimizers. We aim to tackle this with a systematic investigation, which is conducted by analogy with gradient-based model optimizers.
Analogical Analysis
In this section, we present an analogical analysis between model optimization and prompt optimization to build connections and improve existing LLM-based prompt optimizers.
@@ Introduction
Update direction Descent direction Current information
@@
Background
In this part, we first introduce the task definition of LLM-based prompt optimization, and then establish connections with gradient-based optimizers.
Task Formulation.
Prompt optimization aims to find the optimal task prompt $p^{*}$ in the format of natural language that maximizes the performance on a specific task dataset $\mathcal{D}$ when using an LLM as the task model $\mathcal{M}_{T}$. To perform such optimization, our idea is to develop a prompt optimizer, which can be built upon some search algorithm (e.g., evolutionary algorithms (Guo et al. 2023)) or an LLM (Yang et al. 2023). In this paper, we focus on using an LLM as the prompt optimizer $\mathcal{M}_{O}$. Formally, the problem of prompt optimization can be formulated as:
@@
where $p$ is the prompt generated by the prompt optimizer $\mathcal{M}_{O}$, $\mathcal{M}_{T}(x;p)$ represents the output from the task model for input $x$ conditioned on the prompt $p$, and the function $F(\cdot)$ calculates the task performance based on some measurement.
Revisiting Gradient-based Optimizers.
Similar to LLM-based prompt optimizer, gradient-based model optimizer aims to find the optimal values of model parameters that minimize the loss function. In the basic form of gradient descent (Boyd and Vandenberghe 2014), a single optimization step can be formulated as follows:
@@
where $\Theta_{k}$ and $\Theta_{k+1}$ are the values of model parameters at the last and current steps, $\tau_{k}$ and $g_{k}$ are the learning rate and gradient at the current step. Gradient descent can be improved by focusing on two elements in the formula: $\tau_{k}$ and $g_{k}$. For $\tau_{k}$, learning rate schedulers (Gotmare et al. 2019) are proposed to dynamically adjust the learning rate. For $g_{k}$, the concept of momentum (Sutskever et al. 2013) is introduced to include historical gradients, and its computation can be expressed as follows: $v_{k+1}=\beta v_{k}+g_{k}=\sum_{i=1}^{k}\beta^{k-i}g_{i}$, where $\beta$ represents the momentum coefficient.
@@
Despite various gradient-based optimizers, they mainly model two key factors, namely update direction (e.g., gradient $g_{k}$) and update method (e.g., subtract $\tau_{k}g_{k}$). Our approach is inspired by the observation that existing LLM-based prompt optimization methods also implicitly employ the two aspects (see Table 1). However, existing work only initially explores the design of the two key factors, we aim to conduct more in-depth and systematic investigations from these two novel perspectives in the following subsection.
Update Direction
The update direction refers to the adjustments based on the information of previous or current prompts to determine the best direction for improving them. We apply the descent direction and momentum concepts to design the meta-prompts.
Analogical Descent Direction.
Descent direction determines the direction of parameter updates based on the model performance. We analogize two similar forms that determine how to improve new prompts according to the current information.
Prompt Optimizer GPT-3.5-turbo GPT-4 GPT-4
Task Model Llama-2-7b-chat Llama-2-7b-chat GPT-3.5-turbo
Gradient P+M P+M+R P+M P+M+R P+M P+M+R
Momentum None 41.07 40.34 40.32 39.56 64.76 64.55
Summarization 41.03 40.63 40.58 40.41 64.62 64.55
Recency 41.93 41.55 42.02 41.34 65.26 64.97
Relevance 42.53 40.81 42.89 41.97 65.87 65.39
Importance 41.84 39.73 41.73 41.04 65.26 65.11
@@
Table 2: The performance comparison of various update directions based on different prompt optimizers and task models. “P” denotes prompt, “M” denotes performance, and “R” denotes reflection.
$\bullet$ Prompt+Performance. One straightforward method is to include the last-round task prompt and the corresponding model performance into the meta-prompt for the LLM-based optimizer $\mathcal{M}_{O}$. It leverages the capacity of LLMs to reason about how to improve prompting optimization.
$\bullet$ Prompt+Performance+Reflection. Another way to solve the barrier of the descent direction is to leverage the reflection capability of LLMs (Pryzant et al. 2023). With the reflection mechanism, LLMs can generate feedback from past failures to improve performance. Such feedback can be seen as a form of “semantic” gradient signals.
Analogical Momentum.
Inspired by the momentum in gradient descent, we aim to make LLMs aware of the prompts used in the optimization process and their corresponding results (i.e., descent direction), thereby achieving a more global update direction. A straightforward way is to directly include all intermediate results into the meta-prompt. However, it might be limited by the context length of LLMs and affected by the accumulated noise. To better utilize the optimization trajectory, we propose two alternative methods.
$\bullet$ Summarization-based trajectory. One simple approach is to summarize the intermediate results from the optimization trajectory. Specifically, at each step, we use an instruction (i.e., “Your task is to summarize the problems in the previous prompt and the current prompt.”) to let the LLM perform summarization using the summary in the last step and the result in the current step.
$\bullet$ Retrieval-based trajectory. Another approach is to select $k$ pieces of intermediate results from the optimization trajectory. Specifically, we consider three strategies: (1) Recency: selecting $k$ nearest intermediate results; (2) Relevance: selecting $k$ most relevant intermediate results, which are measured by the semantic similarity based on the BGE model (Xiao et al. 2023); (3) Importance: selecting $k$ most important intermediate results, which are measured by the performance.
Update Method
The update method can refer to how the LLM utilizes such information to improve task prompts. Accordingly, we explore how to analogize the learning rate and the descent process into the design of meta-prompts.
Analogical Learning Rate.
In the model optimization, the learning rate controls the extent of gradient updates at each step. Similarly, we aim to control the variation degree of prompt optimization. Specifically, we limit the number of words that can be modified in the meta-prompt (i.e., edit distance). Accordingly, we propose two methods of controlling edit distance as follows:
$\bullet$ Decay-based constraint. To avoid overshooting the optimal solution, the decay strategy is proposed to gradually reduce the learning rate (Izmailov et al. 2018). Here, we borrow the idea of controlling the maximum edit distance and consider gradually reducing its value following either a linear or cosine curve. In particular, we reduce the constraint to approximately 20% of its maximum value until convergence.
$\bullet$ Warmup-based constraint. To avoid instability in the early stage of optimization, the warmup strategy is proposed to gradually increase the learning rate at the beginning (Goyal et al. 2017). Similarly, we adopt the widely used linear warmup schedule to gradually increase the constraint for the maximum edit distance in the initial 5% steps.
Analogical Gradient Descent.
By analogy with the subtraction operation in gradient descent (i.e., $-\tau_{k}g_{k}$ in Eq. (2)), we introduce two methods to update the task prompt accordingly.
$\bullet$ Editing-based refinement. The first method directly edits the last-round task prompt to improve performance. Specifically, we add an instruction (i.e., “Modify the current prompt”) into the meta-prompt, which requires the LLM to edit the task prompt from the last step according to the update direction. This method allows for effectively exploiting the existing prompt.
$\bullet$ Generation-based refinement. In addition to direct edits, we can also leverage the in-context learning capability of LLMs to generate refined task prompts. Specifically, we present the information regarding the updated direction in the format of demonstrations. Then, We add an instruction (i.e., “Write a new improved prompt”) to let the LLM follow the demonstration to directly generate a new task prompt. Compared with the editing-based method, the generation-based approach explores a wider range of prompt variations.
Prompt Optimizer GPT-3.5-turbo GPT-4 GPT-4
Task Model Llama-2-7b-chat Llama-2-7b-chat GPT-3.5-turbo
Learning rate Editing Generation Editing Generation Editing Generation
Descent No control 42.53 42.61 42.89 43.17 65.87 66.37
Fixed 42.91 43.09 43.38 43.66 66.48 66.91
+Warmup 41.76 40.08 42.53 42.95 65.79 65.52
Linear decay 42.68 42.86 43.55 44.03 66.56 67.10
+Warmup 41.47 41.12 41.47 42.91 66.03 66.18
Cosine decay 42.95 43.75 43.98 44.97 66.74 67.80
+Warmup 40.19 41.29 42.68 43.13 65.94 66.37
@@
Table 3: The performance comparison of various update methods based on different prompt optimizers and task models.
Empirical Experiments
In this part, we describe the experiment setting for our analogical analysis and detail our findings from the experiment.
Experiment Setup.
We select a dataset from each type of task in BBH (Suzgun et al. 2023) to create a lite BBH benchmark for our analysis: i) Navigate (binary choice); ii) Movie Recommendation (multiple choice); iii) Object Counting (numeric response); iv) Word Sorting (free response). We adopt exact match as the metric for performance evaluation. We use three different model combinations of prompt optimizers and task models (i.e., GPT-3.5-turbo and Llama-2-7b-chat, GPT-4 and Llama-2-7b-chat, GPT-4 and GPT-3.5-turbo). The optimization process lasts for at most 3 epochs, under which the task prompt can reach the plateau.
Findings for Update Direction.
The results for the update direction are presented in Table 2. Here are the main findings:
$\bullet$ Reflection Leads to Performance Drop. Regarding the analogy to descent direction, prompt+performance achieves better performance than prompt+performance+reflection, which brings an improvement of up to 31% compared with the initial task prompt. The improvement brought by prompts can be attributed to their rich semantic information about the task, which can activate the task-specific knowledge of LLMs for optimization. In contrast, LLMs are limited in their capabilities of reflection (Huang et al. 2023), which may lead to inaccurate updates.
$\bullet$ Prompt Optimizers can Learn More from Contextually Relevant Prompts. The analogical concept of momentum can further improve the performance. Among various designs, relevance-based trajectory emerges as the most effective one, which brings an additional 15% improvement. This can be attributed to the fact that LLMs can learn more from contextually relevant prompts, while it might be challenging for LLMs to fully understand the signal of recency or importance. By contrast, the summarization-based trajectory proves to be less helpful. One possible reason is that summarization only captures common aspects of the trajectory while neglecting details that may be crucial.
Findings for Update Method.
@@
To investigate the update method for prompt optimization, we conduct experiments using the best configuration found in the previous experiments. The results for the update method are presented in Table 3.
@@
$\bullet$ Generation-based Refinement is Better. In general, generation-based refinement outperforms editing-based refinement, which brings an improvement of up to 36% compared with the initial task prompt. This can be partly attributed to the significance of exploration in prompt optimization. Generation-based strategy is not confined to the current prompt, allowing the model to better leverage the in-context examples. Therefore, this approach can demonstrate great flexibility to enable LLMs to explore a larger search space.
@@
$\bullet$ Decay Strategy is Helpful. Among various controlling methods for prompt variation, cosine decay-based constraint achieves the best performance, bringing an additional 10% improvement. However, unlike gradient-based model optimization, the warmup strategy does not yield improvement. This finding suggests that, at the early stage of prompt optimization, exploration plays a crucial role in conducting large-scale prompt searches, while stability becomes more important in the later stage for more refined adjustments.
Method